OpenStudy (anonymous):

The volume of a cube is increasing at a rate of 9 cubic centimetres per second. How fast is the surface area increasing when the length of an edge is 10 centimetres ?

OpenStudy (anonymous):

@bob-bob-bob

OpenStudy (anonymous):

V (cube) = a^3 d v / dt = 3 a^2 d a/ dt

OpenStudy (anonymous):

9 = 3 a^2 d a /dt 3/ a^2 = d a/ dt S = 6 a^2 d s/ dt = 6 * 2 a da / dt

OpenStudy (anonymous):

a = 10.... just sunstistute and find the answer

OpenStudy (anonymous):

Thanks :) @Yahoo!

OpenStudy (anonymous):

Welcome..:)